Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, November 25, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 25, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/25/98 (Wednesday)

Segment 1 report for 11/25/98: Re-seed of Row was under way on Sewer Line Rd. The bore operation continued at Cuddy Creek Tributary (EMP 25.47). On Lebec Rd., coating occurred from EMP 21.48 to 21.49. On Tejon Bench some excavation took place at EMP 18.10. Crews are repairing the Cal Trans fence along I-5. Shading occurred to EMP 17.08. Tie-in of two long segments took place at ~EMP 16.92. On Grapevine Rd. pipe bending and stringing occurred from 15.46 to 15.37. Excavation of the bore pit at Stream Crossing #9 (EMP 15.35) has begun on the north side. Cement pouring occurred at Grapevine Pump Station for the encasement wall around the substation. Pipe fitting and electrical work continued. A crew worked on clean-up along Road B. On Hwy 166, tie-in was completed (except for at Emidio Pump Station). Conduit was placed and back-fill occurred between EMP .53 and .60. At Emidio Pump Station phone line was being installed. Work on the office interior continued.

Segment 2B (Castaic to Sylmar): Report for 11-25-98. Work was limited to backfilling, running 200' of fiber optic conduit and general clean up at Calgrove Blvd and the Old Rd. In addition, crews at Feedmill Rd continued work pulling the sleeve for the directional drill.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 25 November 1998: 3A: The bore crew at Tuxford Street progressed an additional 60 feet today. The tie in of pipe on either side of Tujunga Wash was completed today, with welding, x-ray, and coating crews spending they day on that work. Backfill crews roughly finished the south side of Van Nuys, then relocated to work on the south side of Tujunga was most of the day. Crews at Pacoima Wash finished the placement of the block valve, as well as doing some ditch work . 3C: Ditching crews began work for the tie in to the Hubbard bore. Backfill continued to work between Sepulveda Boulevard and the south side of the California Aqueduct crossing. Welding, x-ray, and coating worked on the tie in to Segment 2, under Balboa overpass. The crew for the cover across from the aqueduct poured more concrete.

Segment 4B (Lynwood to El Segundo): Report for 11-25-98. Crews worked on general clean up of water damage to the right of way after a local fire hydrant was struck by a vehicle. At Lynwood station, crews were working to install the launcher assembly and associated valves. At El Segundo, welders continued work on the valve bodies while concrete was poured for pipe supports. Segment 5: The directional drill under the 91 Freeway was completed today and crews began a general clean up of the drilling site. At the Tosco refinery, boring crews progressed under the roadway east of the plant while welders worked to complete assembly of valve sites.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page